OpenStudy (anonymous):

Solve by the add and subtract method (elimination) x + y = 13 x- y = 7

OpenStudy (anonymous):

solve this by addition of equations add the similar terms x+x=2x y+(-y)=0 13+7=20 then the product equation will be 2x=20 divide both sides by 2 x=10

OpenStudy (anonymous):

What about Y?

OpenStudy (anonymous):

and then plug in the value of x=10 in one of the equations x+y=13 10+y=13 we subtract 10 from both sides y=13-10 y=3

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

@BigBoyLP

OpenStudy (anonymous):

Yes thank you @Abdulhameed

OpenStudy (anonymous):

@BigBoyLP anytime :)